OpenStudy (anonymous):

Write the product in standard form. ( 7 + 7i)( 6 + 7i) a. -7 + 91i b. 49i2 + 91i + 42 c. 91 - 7i d. -7 - 91i

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[(7+7i)(6+7i)=(7)(6)+(7)(7i)+(7i)(6)+(7i)(7i)\\\\ =42+49i+42i+49i^2\\\\=42+91i+49(-1)\\\\ =42-49+91i\\\\\\=-7+91i \]